OpenStudy (anonymous):

Whats does the series from n=0 to infinity 3^(1-n) converge to?

OpenStudy (zarkon):

\[3^{1-n}=\frac{1}{3^{n-1}}=3\frac{1}{3^n}\]

OpenStudy (zarkon):

this is a geometric sum

OpenStudy (anonymous):

Right, and I know the formula for the sum of a geometric series is a/1-r. But I cant get the series into the right form to use that.

OpenStudy (zarkon):

\[\sum_{n=0}^{\infty}3\frac{1}{3^n}=3\sum_{n=0}^{\infty}\left(\frac{1}{3}\right)^n\]

OpenStudy (anonymous):

But dont I need it in the form ar^(n-1) where n greater than or equal to 0?

OpenStudy (zarkon):

that is if the sum starts from n=1

OpenStudy (zarkon):

yours starts at n=0

OpenStudy (zarkon):

you can modify what you have \[\sum_{n=0}^{\infty}3\frac{1}{3^n}=3\sum_{n=0}^{\infty}\left(\frac{1}{3}\right)^n\] \[=3\sum_{n=1}^{\infty}\left(\frac{1}{3}\right)^{n-1}\]

OpenStudy (anonymous):

Ahh I see, so my sum should come to 9/2?

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

Awesome! thanks for the help!

OpenStudy (zarkon):

np